Cull Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on total fixed manufacturing overhead cost of $442,500, variable manufacturing overhead of $2.50 per machine-hour, and 59,000 machine-hours. The company has provided the following data concerning Job X455 which was recently completed:

Number of units in the job 10
Total machine-hours 80
Direct materials $ 870
Direct labor cost $ 1,740

If the company marks up its unit product costs by 20% then the selling price for a unit in Job X455 is closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Fixed manufacturing overhead cost = $442,500

Variable manufacturing overhead cost = $2.50 per machine hour

Machine hours = 59,000

Predetermined overhead rate = Variable manufacturing overhead cost + Fixed manufacturing overhead cost/Machine hours

= 2.50 + 442,500/59,000

= 2.50 + 7.50

= $10 per machine hour

Actual machine hour used by job X455 = 80

Overhead applied to job X455 = Actual machine hour used by job X455 x  Predetermined overhead rate

= 80 x 10

= $800

Job cost sheet (Job X455)

Direct material 870
Direct labor 1,740
Applied overhead 800
Total cost $3,410

Mark up = 20%

= 3,410 x 20%

= $682

Total sales = Cost + mark up

= 3,410 + 682

= $4,092

Selling price per unit = Total sales/Number of units

= 4,092/10

= $409.2

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.